1.write true or false. multiplying each number of a Pythagorean triple be a nonzero whole number yields another Pythagorean triple?
2. The longest side of a triangle is 13 centimeters. another side is 11 centimeters. if the triangle is obtuse. right an inequality for the range of values of the third side?
: 1.write true or false. multiplying each number of a Pythagorean triple be a nonzero whole number yields another Pythagorean triple?
2. The longest side of a triangle is 13 centimeters. another side is 11 centimeters. if the triangle is obtuse. right an inequality for the range of values of the third side?

Answer by mangopeeler07(442)

1.write true or false. multiplying each number of a Pythagorean triple be a nonzero whole number yields another Pythagorean triple?

true.
----------------------------------------------------------------------------
2. The longest side of a triangle is 13 centimeters. another side is 11 centimeters. if the triangle is obtuse. right an inequality for the range of values of the third side?

for a right triangle:
a^2+b^2=c^2
11^2+b^2=13^2
121+b^2=169
b^2=48
b=4sqrt(3)

for an obtuse triangle:
a^2+b^2<c^2
11^2+b^2<13^2
121+b^2<169
b^2<48
b<4sqrt(3)

If it is obtuse, b definitely has to be less than 4sqrt3.

c<a+b
13<11+b
2<b
b also must be more than 2, so that a+b>c and b+c>a. So,

Answer:
2<b<4sqrt(3)